OpenStudy (anonymous):

HELP HELP HELP! IM SO LOST! a farmer has 3000 ft. of fencing available to enclose a rectangular feild. Express the area (A) of the rectangle as a function of X, where X is the length of the rectangle.

OpenStudy (anonymous):

A = W * X where W is the width so 3000 = WX

OpenStudy (anonymous):

gotcha...

OpenStudy (anonymous):

sorry, 3000 is the perimeter. so 3000 = 2(X+W) therefore W = 1500-X so the area is A = WX = X*(1500-X) A = 1500X-X^2

OpenStudy (anonymous):

can u explain a bit more please?

OpenStudy (anonymous):

it is given that he has 3000 ft of fence to fence in the field. that must mean that 3000 feet is the length of the fence around the field. since the field is rectangular, the length of the fence must be the perimeter of the field so 3000 = perimeter of field but for a rectangle, perimeter is 2(length+width) so 3000 = 2(W+X)